1. The boiling point of ethanol, CH3CH2OH, is 78.500 °C at 1 atmosphere. Kb(ethanol) = 1.22 °C/m

In a laboratory experiment, students synthesized a new compound and found that when 11.40 grams of the compound were dissolved in278.0 grams of ethanol, the solution began to boil at 78.673 °C. The compound was also found to be nonvolatile and a non-electrolyte.


What is the molecular weight they determined for this compound ?

____ g/mol

2. Match the following aqueous solutions with the appropriate letter from the column on the right.

1. 0.16 m FeI2 A. Lowest freezing point
2. 0.15 m CoBr2 B. Second lowest freezing point
3. 0.22 m NH4NO3 C. Third lowest freezing point
4. 0.40 m Urea(nonelectrolyte) D. Highest freezing point
Answer #1

1)

Δ Tb = 78.673 - 78.500 = 0.17 oC

use:

Δ Tb = Kb*mb

0.173 = 1.22 *mb

mb= 0.1418 molal

m(solvent)= 278.0 g

= 0.278 kg

use:

number of mol,

n = Molality * mass of solvent in Kg

= (0.1418 mol/Kg)*(0.278 Kg)

= 3.942*10^-2 mol

mass(solute)= 11.40 g

use:

number of mol = mass / molar mass

3.942*10^-2 mol = (11.4 g)/molar mass

molar mass = 289.2 g/mol

Answer: 289.2 g/mol

Only 1 question at a time please